OpenStudy (anonymous):

Divide: 7a+28 over a-4 divided by a2+a-12 over 14a

OpenStudy (pfenn1):

\[\frac{\frac{7a+28}{a-4}}{\frac{a2+a-12}{14a}}\] The first thing I would do is change this expression. Are you aware that dividing one fraction by another is the same as multiplying the top fraction by the inverse of the bottom fraction?

OpenStudy (pfenn1):

\[\frac{\frac{7a+28}{a-4}}{\frac{a^2+a-12}{14a}}=\frac{7a+28}{a-4} \times \frac{14a}{a^2+a-12} \]